I need help?
The sum of two numbers is 33. Their difference is 11. What are the two numbers?

Answer by scott8148(6628)

x+y=33 ... x-y=11 ... adding the equations gives (x+y)+(x-y)=33+11 ... 2x=44 ... x=22

(22)+y=33 ... y=11